# GridRoute: A Benchmark for LLM-Based Route Planning with Cardinal Movement in Grid Environments

### v1

@@ A.3 AoT with Example (Algo-Shot) Prompt
"Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use Dijkstra’s algorithm to calculate the shortest path. Follow these steps:" 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Dijkstra’s algorithm: - Assign an initial distance of infinity to all grid cells except the starting point, which should have a distance of 0. - Use a priority queue to repeatedly select the grid cell with the smallest tentative distance. - For each selected cell, evaluate its valid neighbors and update their tentative distances if a shorter path is found. - Continue until the endpoint ({end_x}, {end_y}) is reached or all possible paths have been explored. 5. Construct the path:" - Start at ({start_x}, {start_y}). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Example 1: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))] Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (5, 3), (4, 3)] Example 2: Start: (2, 4), End: (7, 5), Obstacles: [((3, 4), (5, 6))] Path: [(2, 4), (2, 3), (3, 3), (4, 3), (5, 3), (6, 3), (7, 3), (7, 4), (7, 5)] Return only the path as output, without additional explanations.
@@ Appendix B Results
| deepseek-V3 | AoT | 100 | 79 | 65 | 111.64 | 5.87 | 8.38 | 100 | 67 | 62 | 105.28 | 25.43 | 10.26 | 100 | 72 | 59 | 114.34 | 131.77 | 14.40 |
| qwen2.5-7b-instruct | (DFS) | 100 | 51 | 47 | 105.68 | 2.11 | 2.59 | 100 | 27 | 26 | 101.19 | 2.37 | 3.88 | 100 | 19 | 15 | 106.26 | 9.89 | 12.39 |
@@
| deepseek-V3 | AoT | 100 | 80 | 80 | 100 | 0 | 3.27 | 100 | 74 | 74 | 100 | 0 | 7.13 | 100 | 66 | 66 | 100 | 0 | 6.43 |
| qwen2.5-7b-instruct | (A*) | 100 | 57 | 55 | 101.16 | 1.19 | 3.03 | 100 | 36 | 34 | 102.97 | 8 | 7.85 | 99 | 25 | 24 | 104.49 | 36 | 7.13 |
@@
| deepseek-V3 | AoT | 100 | 81 | 79 | 101.40 | 0.49 | 1.89 | 100 | 69 | 69 | 100 | 0 | 2.57 | 100 | 73 | 72 | 101.13 | 17.75 | 3.46 |
| qwen2.5-7b-instruct | (Dijkstra) | 100 | 57 | 55 | 1.0094 | 0.35 | 2.00 | 100 | 33 | 31 | 101.88 | 4.48 | 2.87 | 100 | 23 | 20 | 106.38 | 17.04 | 66.42 |

### v2

@@ A.3 AoP with Example (Algo-Shot) Prompt
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use Dijkstra’s algorithm to calculate the shortest path. Follow these steps:" 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Up: (x, y+1) - Down: (x, y-1) - Right: (x+1, y) - Left: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Dijkstra’s algorithm: - Assign an initial distance of infinity to all grid cells except the starting point, which should have a distance of 0. - Use a priority queue to repeatedly select the grid cell with the smallest tentative distance. - For each selected cell, evaluate its valid neighbors and update their tentative distances if a shorter path is found. - Continue until the endpoint ({end_x}, {end_y}) is reached or all possible paths have been explored. 5. Construct the path:" - Start at ({start_x}, {start_y}). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Example 1: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))] Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (5, 3), (4, 3)] Example 2: Start: (2, 4), End: (7, 5), Obstacles: [((3, 4), (5, 6))] Path: [(2, 4), (2, 3), (3, 3), (4, 3), (5, 3), (6, 3), (7, 3), (7, 4), (7, 5)] Return only the path as output, without additional explanations.
@@ Appendix B Results
| deepseek-V3 | AoP | 100 | 79 | 65 | 111.64 | 5.87 | 8.38 | 100 | 67 | 62 | 105.28 | 25.43 | 10.26 | 100 | 72 | 59 | 114.34 | 131.77 | 14.40 |
| qwen2.5-7b-instruct | DFS | 100 | 51 | 47 | 105.68 | 2.11 | 2.59 | 100 | 27 | 26 | 101.19 | 2.37 | 3.88 | 100 | 19 | 15 | 106.26 | 9.89 | 12.39 |
@@
| deepseek-V3 | AoP | 100 | 80 | 80 | 100 | 0 | 3.27 | 100 | 74 | 74 | 100 | 0 | 7.13 | 100 | 66 | 66 | 100 | 0 | 6.43 |
| qwen2.5-7b-instruct | A* | 100 | 57 | 55 | 101.16 | 1.19 | 3.03 | 100 | 36 | 34 | 102.97 | 8 | 7.85 | 99 | 25 | 24 | 104.49 | 36 | 7.13 |
@@
| deepseek-V3 | AoP | 100 | 81 | 79 | 101.40 | 0.49 | 1.89 | 100 | 69 | 69 | 100 | 0 | 2.57 | 100 | 73 | 72 | 101.13 | 17.75 | 3.46 |
| qwen2.5-7b-instruct | Dijkstra | 100 | 57 | 55 | 1.0094 | 0.35 | 2.00 | 100 | 33 | 31 | 101.88 | 4.48 | 2.87 | 100 | 23 | 20 | 106.38 | 17.04 | 66.42 |